# DanmakuTPPBench: A Multi-modal Benchmark for Temporal Point Process Modeling and Understanding

### v1

@@ Abstract
We introduce DanmakuTPPBench, a comprehensive benchmark designed to advance multi-modal Temporal Point Process (TPP) modeling in the era of Large Language Models (LLMs). While TPPs have been widely studied for modeling temporal event sequences, existing datasets are predominantly unimodal, hindering progress in models that require joint reasoning over temporal, textual, and visual information. To address this gap, DanmakuTPPBench comprises two complementary components: (1) DanmakuTPP-Events, a novel dataset derived from the Bilibili video platform, where user-generated bullet comments (Danmaku) naturally form multi-modal events annotated with precise timestamps, rich textual content, and corresponding video frames; (2) DanmakuTPP-QA, a challenging question-answering dataset constructed via a novel multi-agent pipeline powered by state-of-the-art LLMs and multi-modal LLMs (MLLMs), targeting complex temporal-textual-visual reasoning. We conduct extensive evaluations using both classical TPP models and recent MLLMs, revealing significant performance gaps and limitations in current methods’ ability to model multi-modal event dynamics. Our benchmark establishes strong baselines and calls for further integration of TPP modeling into the multi-modal language modeling landscape. The code and dataset have been released at https://github.com/FRENKIE-CHIANG/DanmakuTPPBench
@@ 1 Introduction
Temporal Point Processes (TPPs) offer a powerful framework for modeling event sequences in continuous time and have shown effectiveness across a wide range of domains, including social media activity prediction, healthcare monitoring, earthquake modeling, and financial transaction analysis [18, 11, 4, 6, 9, 2]. However, conventional TPP models often struggle to capture the rich multi-modal patterns and contextual dependencies present in real-world event streams. Meanwhile, Large Language Models (LLMs) and Multi-modal LLMs (MLLMs) have recently achieved impressive success across various tasks [21, 1]. Integrating TPP modeling into these architectures presents a promising yet largely underexplored direction, with the potential to enhance temporal reasoning capabilities and unlock novel downstream applications [16, 13].
@@
We first introduce DanmakuTPP-Events, a novel multi-modal TPP dataset constructed from Bilibili’s Danmaku (real-time bullet comment) system. Danmaku refers to user-generated comments that appear as overlaid text at specific timestamps during video playback, creating an interactive and communal viewing experience. As one of China’s leading video-sharing platforms, Bilibili offers an ideal source for TPP data—each Danmaku comment naturally forms an event with precise temporal alignment, rich textual content, and corresponding visual context from video frames. This makes it a native setting for modeling temporal, textual, and visual modalities in combination. The resulting dataset contains 7,250 TPP sequences and over 10.8 million Danmaku events.
@@ 2 Related Works
We summarize widely used datasets for Temporal Point Process (TPP) modeling in Table 1. These datasets primarily focus on capturing temporal patterns and event types across various application domains. The Retweet dataset [34] contains user retweet sequences, categorized into three groups based on follower counts. The StackOverflow dataset [10] records user badge-awarding events with 22 distinct badge types. The Taobao dataset [26] captures user click behavior across 17 item categories, while the Taxi dataset [25] logs pick-up and drop-off events in New York City, categorized into 10 location-based event types. Although these datasets provide valuable temporal signals, they lack multi-modal context. The ActiveRT dataset [22] contains tweets linking to YouTube videos collected over six months, introducing some content-level variation. The RNCNIX[12] and Amazon Review[19] datasets incorporate textual information—news articles and customer reviews, respectively—offering partial multi-modality. However, these datasets still lack visual components, limiting their applicability for modeling richly contextualized event sequences. In contrast, our proposed DanmakuTPP-Events dataset is the first to natively integrate temporal, textual, and visual modalities by leveraging timestamped user comments overlaid on video content. This enables more comprehensive multi-modal TPP modeling in a naturally aligned setting.
Table 1: Comparison between prior TPP datasets and our proposed DanmakuTPPBench. K denotes the number of event types, $L_{avg}$ stands for the average TPP sequence length.
@@
DanmakuTPP-Events 7,250 10 1494
DanmakuTPP-QA 2,605 10 967 (10 tasks)
@@ 3.1 Construction of DanmakuTPP-Events Dataset
Bilibili is a leading Chinese video-sharing platform where users can upload, watch, and annotate videos with Danmaku comments. We manually collect all videos posted by the top 100 most popular creators on Bilibili in 2024111The full list of creators is available at https://www.bilibili.com/BPU2024., resulting in a total of 7,250 videos. From these videos, we construct a large-scale multi-modal TPP dataset, DanmakuTPP-Events, which includes 10,820,790 Danmaku events. Formally, for a given video $v$, the Danmaku sequence is represented as $\operatorname{\mathcal{H}}_{v}={(t_{i},e_{i},m^{\text{text}}_{i},m^{\text{image}}i)}{i=1}^{N}$, where each event $i$ comprises a timestamp $t_{i}$, event type $e_{i}$, a textual mark $m^{\text{text}}_{i}$, and an associated video frame $m^{\text{image}}_{i}$. We preprocess and format the dataset following the protocol established in [27].
Data Statistics. As shown in Fig. 2 (a), DanmakuTPP-Events spans fourteen distinct video categories. Gaming constitutes the largest share (23%), followed by education (18%), lifestyle (12%), and animation (10%). This diverse topical distribution ensures broad domain coverage, enhancing the dataset’s ecological validity for TPP modeling.
Fig. 2 (b) presents the distribution of video durations, revealing that most videos fall within the 0–500 second range. The distribution exhibits a long-tail pattern, with a small number of videos exceeding 1000 seconds and a few reaching up to 5000 seconds. These variations introduce significant diversity in the temporal structure of Danmaku activity.
Fig. 2 (c) shows the distribution of Danmaku comment counts per video. While the majority of videos receive 500–2000 comments, there are pronounced peaks around 1000 and 3500, with some videos attracting 8000–10000 comments. This wide variation in comment density gives rise to a rich variety of temporal patterns, ranging from dense comment bursts to more dispersed interactions over time.
Together, the heterogeneity in video topics, durations, and comment densities contributes to the complexity and expressiveness of the DanmakuTPP-Events dataset, making it a valuable resource for advancing multi-modal TPP modeling.
@@ 3.2 Multi-Agent Collaboration for DanmakuTPP-QA Construction
To enable temporal-textual-visual reasoning, we curated a subset of 2,605 videos from the DanmakuTPP-Events dataset, each containing between 500 and 1,500 Danmaku events. Based on this selection, we developed a multi-agent collaborative framework for automated data construction, resulting in a challenging multi-task benchmark: DanmakuTPP-QA. As shown in Fig. 3, our pipeline integrates five specialized agents, each responsible for a key stage in the data generation process:
@@
(4) Visualization Agent. This agent transforms temporal sequences and annotated attributes into interpretable visualizations. Using Qwen2.5-Coder [8], Python scripts are automatically generated to produce charts such as Danmaku intensity timelines, sentiment evolution plots, and event-type distributions. These visualizations enrich the raw data with aligned textual-temporal-visual representations.
@@
where $<|TPP\_data|>$ serves as a placeholder for the corresponding TPP data associated with each video. Detailed prompt designs for all agents are provided in Appendix A. To ensure evaluation accuracy, all answers generated by the Task-solve Agent group for the test set underwent rigorous manual verification. This process guarantees the high fidelity of the ground-truth annotations, establishing a reliable foundation for downstream model evaluation and performance benchmarking.
Composition. DanmakuTPP-QA is divided into three subsets: a training set (2,005 samples), a validation set (300 samples), and a test set (300 samples). As summarized in Table 2, the benchmark comprises 10 distinct tasks, including 8 TPP-focused QA tasks and 2 higher-level temporal-textual-visual reasoning tasks. These tasks collectively span a wide range of challenges, such as predicting the timing and burst peaks of Danmaku events, classifying their trigger types, analyzing sentiment trends, and identifying the underlying causes of Danmaku bursts.
@@ 4.1 Experimental Settings
DanmakuTPP-Events Evaluation. We employ the following state-of-the-art deep TPP models when evaluating TPP tasks on DanmakuTPP: (1) Neural Hawkes Process (NHP) [17]: NHP employs a continuous-time LSTM to encode the temporal and type information of historical events. The resulting history embedding is used to model the conditional intensity function. (2) Self-Attentive Hawkes Process (SAHP) [31], Transformer Hawkes Process (THP) [35] and Attentive neural Hawkes proces (AttNHP) [30]: all three models encode historical events through self-attention mechanism. We use the implementation provided by EasyTPP [27]. All models are trained with a batch size of 64 for 100 epochs using the Adam optimizer with a learning rate of 1e-3 and early stopping based on validation performance. We note that the multi-modal information is not used for these models.
DanmakuTPP-QA Evaluation. We frist evaluate the TPP understanding capabilities of existing open source pretrained LLMs (Qwen2.5 series [28], Qwen3 series [29] and Deepseek-V3 [15]) and MLLMs (Gemma3 [23] and Qwen2.5-VL [24]). According to the different video clips required by each task, we provide the textual timestamps and Danmaku content within the time window as inputs to both LLMs and MLLMs. Additionally, we randomly sample three video frames within the time window for the supplementary input of MLLMs, enhancing the multi-modal understanding. LLMs focus on mining temporal evolution patterns from timestamps and Danmaku, while MLLMs are responsible for aligning textual TPP information with video frames to address complex TPP tasks.
We then further evaluate the model performance after finetuning. We conduct LoRA [7] finetuning with Qwen2.5-VL-3B. For each task, we train the model on a single NVIDIA RTX 4090 for 3 epochs with learning rate 1e-4. Due to GPU memory constraints, we set a truncation of the sequence for some tasks during training. Details about hyperparameters can be found in Appendix B.
Evaluation Metrics. For conventional TPP tasks on DanmakuTPP-Events, we employ RMSE for evaluating next-event timestamp prediction and test log-likelihood values for measuring modeling performance [27]. For DanmakuTPP-QA, we use different metrics to comprehensively assess model performance across various task settings, as shown in Table 2. Tasks related to temporal prediction and sentiment polarity computation adopt root mean square error (RMSE). Tasks related to Danmaku event type prediction are evaluated by accuracy. Open-ended tasks are evaluated using Qwen3-235B-A22 to output a correctness score between 0-1 when compared against ground-truth descriptions.
@@ 4.2 Results on DanmakuTPP-Events
As shown in Fig. 4(a), we observe varying performance across different TPP models on DanmakuTPP-Events dataset. For log-likelihood evaluation, NHP achieves the best performance with a score of 0.755, followed by SAHP (0.704), while THP (0.470) and AttNHP (0.404) show comparatively lower performance. This suggests that the continuous-time LSTM approach in NHP may be more effective at capturing the complex temporal dynamics in Danmaku data than some attention-based approaches.
For next-event prediction evaluated by RMSE where the lower is better, THP performs best with a score of 0.917, followed closely by NHP (0.934) and AttNHP (0.945). SAHP shows the highest error with an RMSE of 1.040. This reveals a performance discrepancy between likelihood modeling and prediction tasks—models that excel at capturing the underlying distribution (high log-likelihood) do not necessarily perform best at point prediction tasks.
@@
4.3 Results on DanmakuTPP-QA
Table 3 presents a comprehensive evaluation of various LLMs and MLLMs on the DanmakuTPP-QA dataset across eight closed-ended TPP tasks. The tasks range from temporal prediction (T-1 to T-3) to sentiment analysis (T-4 to T-6) and event type prediction (T-7 to T-8), collectively assessing models’ capabilities in understanding and reasoning about multimodal temporal point processes. The results reveal several important insights about model performance on TPP understanding tasks.
@@ 4.3 Results on DanmakuTPP-QA
Open-ended TPP Question Performance. Beyond closed-ended tasks, we further evaluate models on open-ended TPP questions from Tasks 9 and 10, which require deeper reasoning about temporal dynamics. As shown in Fig. 4(b), model performance varies significantly across these challenging tasks. For Task-9 (analysis of global sentiment dynamics), Qwen2.5-VL-72B and Qwen3-235B-A22B demonstrate the strongest performance with correctness scores of 0.48 and 0.42, respectively. In Task-10 (causal attribution analysis), Qwen3-235B-A22B achieves the highest score (0.52), substantially outperforming other models. Notably, we observe that while larger models generally perform better, model architecture and pretraining strategies also play crucial roles, as evidenced by DeepSeek-V3’s competitive performance (0.40 on Task-10) despite its smaller parameter count compared to some other models. The finetuned Qwen2.5-VL-3B-Instruct model (finetuned specifically on DanmakuTPP-QA) demonstrates decent performance on both open-ended tasks (0.36 and 0.30 on Tasks 9 and 10, respectively), outperforming several larger pretrained models. This further confirms that targeted finetuning on TPP data can effectively enhance models’ ability to reason about complex temporal patterns, even with relatively limited parameter counts.

### v2

@@ Abstract
We introduce DanmakuTPPBench, a comprehensive benchmark designed to advance multi-modal Temporal Point Process (TPP) modeling in the era of Large Language Models (LLMs). While TPPs have been widely studied for modeling temporal event sequences, existing datasets are predominantly unimodal, hindering progress in models that require joint reasoning over temporal, textual, and visual information. To address this gap, DanmakuTPPBench comprises two complementary components: (1) DanmakuTPP-Events, a novel dataset derived from the Bilibili video platform, where user-generated bullet comments (Danmaku) naturally form multi-modal events annotated with precise timestamps, rich textual content, and corresponding video frames; (2) DanmakuTPP-QA, a challenging question-answering dataset constructed via a novel multi-agent pipeline powered by state-of-the-art LLMs and multi-modal LLMs (MLLMs), targeting complex temporal-textual-visual reasoning. We conduct extensive evaluations using both classical TPP models and recent MLLMs, revealing significant performance gaps and limitations in current methods’ ability to model multi-modal event dynamics. Our benchmark establishes strong baselines and calls for further integration of TPP modeling into the multi-modal language modeling landscape. Project page: https://github.com/FRENKIE-CHIANG/DanmakuTPPBench.
@@ 1 Introduction
Temporal Point Processes (TPPs) offer a powerful framework for modeling event sequences in continuous time and have shown effectiveness across a wide range of domains, including social media activity prediction, healthcare monitoring, earthquake modeling, and financial transaction analysis [21, 14, 6, 8, 12, 2]. However, conventional TPP models often struggle to capture the rich multi-modal patterns and contextual dependencies present in real-world event streams. Meanwhile, Large Language Models (LLMs) and Multi-modal LLMs (MLLMs) have recently achieved impressive success across various tasks [24, 1, 11]. Integrating TPP modeling into these architectures presents a promising yet largely underexplored direction, with the potential to enhance temporal reasoning capabilities and unlock novel downstream applications [19, 16, 26].
@@
We first introduce DanmakuTPP-Events, a novel multi-modal TPP dataset constructed from Bilibili’s Danmaku (real-time bullet comment) system. Danmaku refers to user-generated comments that appear as overlaid text at specific timestamps during video playback, creating an interactive and communal viewing experience, as illustrated in Fig. 1 (a). As one of China’s leading video-sharing platforms, Bilibili offers an ideal source for TPP data—each Danmaku comment naturally forms an event with precise temporal alignment, rich textual content, and visual context from video frames. This makes it a native setting for modeling temporal, textual, and visual modalities in combination. The resulting dataset contains 7,250 TPP sequences and over 10.8 million Danmaku events.
@@ 2 Related Works
We summarize widely used datasets for TPP modeling in Table˜1. These datasets primarily focus on capturing temporal patterns and event types across various application domains. The Retweet dataset [38] contains user retweet sequences, categorized into three groups based on follower counts. The StackOverflow dataset [13] records user badge-awarding events with 22 distinct badge types. The Taobao dataset [30] captures user click behavior across 17 item categories, while the Taxi dataset [29] logs pick-up and drop-off events in New York City, categorized into 10 location-based event types. Although these datasets provide valuable temporal signals, they lack multi-modal context. The ActiveRT dataset [25] contains tweets linking to YouTube videos collected over six months, introducing some content-level variation. The RNCNIX[15] and Amazon Review[22] datasets incorporate textual information—news articles and customer reviews, respectively—offering partial multi-modality. However, these datasets still lack visual components, limiting their applicability for modeling richly contextualized event sequences. In contrast, our proposed DanmakuTPP-Events dataset is the first to natively integrate temporal, textual, and visual modalities by leveraging timestamped user comments overlaid on video content. This enables more comprehensive multi-modal TPP modeling in a naturally aligned setting.
Table 1: Comparison between prior TPP datasets and our proposed DanmakuTPPBench. K denotes the number of event types, $L_{avg}$ stands for the average TPP sequence length. The Danmaku event types identified in our dataset are illustrated in Fig. 1 (b)
@@
DanmakuTPP-Events 7,250 9 1494
DanmakuTPP-QA 2,605 9 967 (10 tasks)
@@ 3.1 Construction of DanmakuTPP-Events Dataset
Bilibili is a leading Chinese video-sharing platform where users can upload, watch, and annotate videos with Danmaku comments. We manually collect all videos posted by the top 100 most popular creators on Bilibili in 2024111The full list of creators is available at https://www.bilibili.com/BPU2024., resulting in a total of 7,250 videos. From these videos, we construct a large-scale multi-modal TPP dataset, DanmakuTPP-Events, which includes 10,820,790 Danmaku events. Formally, for a given video $v$, the Danmaku sequence is represented as $\operatorname{\mathcal{H}}_{v}={(t_{i},e_{i},m^{\text{text}}_{i},m^{\text{image}}_{i})}_{i=1}^{N}$, where each event $i$ comprises a timestamp $t_{i}$, event type $e_{i}$, a textual mark $m^{\text{text}}_{i}$, and an associated video frame $m^{\text{image}}_{i}$. We preprocess and format the dataset following the protocol established in [31].
Ethical Issues. During the data collection process, we strictly adhered to Bilibili’s Terms of Service and only collected data from publicly available videos that did not require any login or authentication. Furthermore, regarding data privacy and user anonymization, all Danmaku comments in the interactive videos are entirely dissociated from any identifiable user information. The dataset we constructed contains only Danmaku text, temporal sequences, and video frames, and excludes any user account details, personally identifiable information (PII), or other privacy-sensitive content. Finally, to mitigate potential harassment or harmful content that may exist in raw data from the video platform, we proactively implemented multiple methodological precautions to ensure ethical compliance, as detailed in the Appendix A.
Data Statistics. As shown in Fig.˜2 (a), DanmakuTPP-Events spans fourteen distinct video categories. Gaming constitutes the largest share (23%), followed by education (18%), lifestyle (12%), and animation (10%). This diverse topical distribution ensures broad domain coverage, enhancing the dataset’s ecological validity for TPP modeling.
Fig.˜2 (b) presents the distribution of video durations, revealing that most videos fall within the 0–500 second range. The distribution exhibits a long-tail pattern, with a small number of videos exceeding 1000 seconds and a few reaching up to 5000 seconds. These variations introduce significant diversity in the temporal structure of Danmaku activity.
Fig.˜2 (c) shows the distribution of Danmaku comment counts per video. While the majority of videos receive 500–2000 comments, there are pronounced peaks around 1000 and 3500, with some videos attracting 8000–10000 comments. This wide variation in comment density gives rise to a rich variety of temporal patterns, ranging from dense comment bursts to more dispersed interactions over time.
Together, the heterogeneity in video topics, durations, and comment densities contributes to the complexity and expressiveness of the DanmakuTPP-Events dataset, offering not just breadth of content, but depth of interaction, and making it a valuable resource for advancing multi-modal TPP modeling.
@@ 3.2 Multi-Agent Collaboration for DanmakuTPP-QA Construction
To enable temporal-textual-visual reasoning, we curated a subset of 2,605 videos from the DanmakuTPP-Events dataset, each containing between 500 and 1,500 Danmaku events. Based on this selection, we developed a multi-agent collaborative framework for automated data construction, resulting in a challenging multi-task benchmark: DanmakuTPP-QA. As shown in Fig.˜3, our pipeline integrates five specialized agents, each responsible for a key stage in the data generation process:
@@
(4) Visualization Agent. It transforms temporal sequences and annotated attributes into interpretable visualizations. Using Qwen2.5-Coder [10], Python scripts are automatically generated to produce charts such as Danmaku intensity timelines, sentiment evolution plots, and event-type distributions. These visualizations enrich the raw data with aligned textual-temporal-visual representations.
@@
where <|TPP_data|><|TPP\_data|> serves as a placeholder for the corresponding TPP data associated with each video. Detailed prompt designs for all agents are provided in Appendix B. To ensure evaluation accuracy, all answers generated by the Task-solve Agent group for the test set underwent rigorous manual verification. This process guarantees the high fidelity of the ground-truth annotations, establishing a reliable foundation for downstream model evaluation and performance benchmarking.
Composition. DanmakuTPP-QA is divided into three subsets: a training set (2,005 samples), a validation set (300 samples), and a test set (300 samples). As summarized in Table˜2, the benchmark comprises 10 distinct tasks, including 8 TPP-focused QA tasks and 2 higher-level temporal-textual-visual reasoning tasks. These tasks collectively span a wide range of challenges, such as predicting the timing and burst peaks of Danmaku events, classifying their trigger types, analyzing sentiment trends, and identifying the underlying causes of Danmaku bursts.
@@ 4.1 Experimental Settings
DanmakuTPP-Events Evaluation. We employ the following state-of-the-art deep TPP models when evaluating TPP tasks on DanmakuTPP: (1) Neural Hawkes Process (NHP) [20] and Recurrent Marked Temporal Point Process (RMTPP) [5] are two RNN-based models, encoding the temporal and type information of historical events. As for NHP, the resulting history embedding is used to model the conditional intensity function. (2) Self-Attentive Hawkes Process (SAHP) [35], Transformer Hawkes Process (THP) [39] and Attentive neural Hawkes proces (AttNHP) [34]: all three models encode historical events through self-attention mechanism. We use the implementation provided by EasyTPP [31]. (3) ODE-based TPP (ODETPP) [4] models the continuous-time evolution of a latent state via a neural ordinary differential equation. It can be viewed as a spatio-temporal point process restricted to the temporal dimension, with the spatial component omitted. All models are trained with the default setting in the EasyTPP framework to ensure a fair and standardized comparison. We note that the multi-modal information is not used for these models.
DanmakuTPP-QA Evaluation. We frist evaluate the TPP understanding capabilities of existing open source pretrained LLMs (Qwen2.5 series [32], Qwen3 series [33] and Deepseek-V3 [18]) and MLLMs (Gemma3 [27] and Qwen2.5-VL [28]). According to the different video clips required by each task, we provide the textual timestamps and Danmaku content within the time window as inputs to both LLMs and MLLMs. Inspired by the mainstream paradigm for video understanding employed by several powerful MLLMs, we process a video as a sequence input of discrete image frames. We randomly sample three video frames within the time window for the supplementary input of MLLMs, enhancing the multi-modal understanding. The strategy of using three frames was intended as an efficient starting point, balancing the need for visual context against the significant computational and memory demands associated with processing long TPP sequences combined with numerous video frames. Consequently, LLMs focus on mining temporal evolution patterns from timestamps and Danmaku, while MLLMs are responsible for aligning textual TPP information with video frames to address complex TPP tasks.
We then further evaluate the model performance after finetuning. We conduct LoRA [9] finetuning with Qwen2.5-VL-3B. For each task, we train the model on a single NVIDIA RTX 4090 for 3 epochs with learning rate 1e-4. Due to GPU memory constraints, we set a truncation of the sequence for some tasks during training. Details about hyperparameters can be found in Appendix C.
Evaluation Metrics. For conventional TPP tasks on DanmakuTPP-Events, we employ RMSE for evaluating next-event timestamp prediction and test log-likelihood values for measuring modeling performance [31]. For DanmakuTPP-QA, we use different metrics to comprehensively assess model performance across various task settings, as shown in Table˜2. Tasks related to temporal prediction and sentiment polarity computation adopt root mean square error (RMSE). Tasks related to Danmaku event type prediction are evaluated by accuracy. Open-ended tasks are evaluated using Qwen3-235B-A22 to output a correctness score between 0-1 when compared against ground-truth descriptions.
@@ 4.2 Results on DanmakuTPP-Events
As shown in Fig.˜4(a), we observe varying performance across different TPP models on DanmakuTPP-Events dataset. For log-likelihood evaluation, NHP achieves the best performance with a score of 0.799, followed by ODETPP (0.729), SAHP (0.713), and RMTPP (0.646), while THP (0.619) and AttNHP (0.550) show comparatively lower performance. This suggests that the continuous-time LSTM approach in NHP may be more effective at capturing the complex temporal dynamics in Danmaku data than some attention-based approaches.
For next-event prediction evaluated by RMSE where the lower is better, NHP and THP perform best with scores of 0.932 and 0.938, respectively, followed closely by ODETPP (0.968) and AttNHP (0.969). SAHP and RMTPP show the highest error with RMSE scores of 1.025 and 1.009, respectively. Taken together, while NHP performs best on both metrics, discrepancies for models like SAHP and THP suggest that modeling the event-time distribution and optimizing next-event accuracy emphasize different inductive biases, making dual-metric evaluation essential.
4.3 Results on DanmakuTPP-QA
Table˜3 presents a comprehensive evaluation of various LLMs and MLLMs on the DanmakuTPP-QA dataset across eight closed-ended TPP tasks. The tasks range from temporal prediction (T-1 to T-3) to sentiment analysis (T-4 to T-6) and event type prediction (T-7 to T-8), collectively assessing models’ capabilities in understanding and reasoning about multimodal temporal point processes. The results reveal several important insights about model performance on TPP understanding tasks.
@@ 4.3 Results on DanmakuTPP-QA
Open-ended TPP Question Performance. Beyond closed-ended tasks, we further evaluate models on open-ended TPP questions from Tasks 9 and 10, which require deeper reasoning about temporal dynamics. As shown in Fig.˜4(b), model performance varies significantly across these challenging tasks. For Task-9 (analysis of global sentiment dynamics), Qwen2.5-VL-72B and Qwen3-235B-A22B demonstrate the strongest performance with correctness scores of 0.48 and 0.42, respectively. In Task-10 (causal attribution analysis), Qwen3-235B-A22B achieves the highest score (0.52), substantially outperforming other models. Notably, we observe that while larger models generally perform better, model architecture and pretraining strategies also play crucial roles, as evidenced by DeepSeek-V3’s competitive performance (0.40 on Task-10) despite its smaller parameter count compared to some other models. The finetuned Qwen2.5-VL-3B-Instruct model (finetuned specifically on DanmakuTPP-QA) demonstrates decent performance on both open-ended tasks (0.36 and 0.30 on Tasks 9 and 10, respectively), outperforming several larger pretrained models. This further confirms that targeted finetuning on TPP data can effectively enhance models’ ability to reason about complex temporal patterns, even with relatively limited parameter counts.